OpenStudy (anonymous):

Yet another Calculus question im having trouble with . Find the particular solution to y " = sin(x) given the general solution y = −sin(x) + Ax + B and the initial conditions y(pi/2) = 0 and y '(pi/2) = -2.

OpenStudy (anonymous):

not even sure where to start on this one

OpenStudy (freckles):

you make a guess for the particular solution I know the derivative of sin(x) is cos(x) and the derivative of cos(x) is -sin(x) so I might try something like \[ y_p=A sin(x)+Bcos(x)\] for the particular solution

OpenStudy (freckles):

you can call A and B something else since they already being used

OpenStudy (anonymous):

I'm given the answer choices do you think they might help?

OpenStudy (usukidoll):

method of undetermined coefficients...

OpenStudy (usukidoll):

for a second order ode....

OpenStudy (anonymous):

i'v never been thought second order.... are they much different?

OpenStudy (freckles):

yea you will get something like -Asin(x)-Bcos(x)=sin(x) shich means B=0 and -A=1 so A=-1 so we didn't need the cos part of the particular solution

OpenStudy (usukidoll):

second order odes are easier than first order odes... that's what a ton of ode students say

OpenStudy (anonymous):

these area the answer choices by the way −sin(x) +1 + π −sin(x) − 2x + π −sin(x) − 2x − π −sin(x) − 2x + π + 1

OpenStudy (freckles):

it looks like it already gave you the solution

OpenStudy (freckles):

just apply the conditions

OpenStudy (anonymous):

what do you mean .... im not even sure what to solve for

OpenStudy (freckles):

\[y=-\sin(x)+Ax+B \\ y(\frac{\pi}{2})=0 \text{ insert this into the solution }\] you will also need to find y' given that and then use the other condition y'(pi/2)=-2

OpenStudy (freckles):

you will wind up with a system of equations

OpenStudy (anonymous):

so you mean \[y(\frac{ \pi }{ 2 }) = -1 + Api + B\]

OpenStudy (freckles):

not exactly \[y(\frac{\pi}{2})=-\sin(\frac{\pi}{2})+A(\frac{\pi}{2})+B \\ 0=-1+A(\frac{\pi}{2})+B\]

OpenStudy (usukidoll):

x will be pi/2 and then that would be = 0 just like freckies put it

OpenStudy (freckles):

now what is y'

OpenStudy (anonymous):

do we have to derive -sin(x) + Ax + B before plugging in pi/2?

OpenStudy (freckles):

yes otherwise you are differentiating constants on both sides and wind up with 0=0 which doesn't really help \[y=-\sin(x)+Ax+B \\ y'=-\cos(x)+A+0 \\ y'=-\cos(x)+A\]

OpenStudy (freckles):

now plug in the condition

OpenStudy (freckles):

for y'

OpenStudy (anonymous):

yay i found a !!!!!!

OpenStudy (freckles):

go back to the first equation and find B

OpenStudy (anonymous):

okay ......aahhh im so exited for some reason !!!

OpenStudy (freckles):

because this was easier than you thought imaginable ?

OpenStudy (anonymous):

possibly ... okay so im getting B = -pi +1 is that right?

OpenStudy (freckles):

hmmm... I might be geting pi+1

OpenStudy (freckles):

one sec

OpenStudy (anonymous):

no yea your right its -2 not 2 my bad

OpenStudy (anonymous):

okay so what now

OpenStudy (freckles):

\[y'(\frac{\pi}{2})=-\cos(\frac{\pi}{2})+A \\ -2=-\cos(\frac{\pi}{2})+A \\ -2 =0+A \\ -2=A \\ 0=-1+A(\frac{\pi}{2})+B \\ 0=-1-2(\frac{\pi}{2})+B \\ 0=-1-\pi+B \\ B=\pi+1\] just wanted to make sure

OpenStudy (freckles):

well as said you already had the solution the conditions just needed to be applied so you found A and B

OpenStudy (freckles):

just plug them back in

OpenStudy (freckles):

into the solution they gave you

OpenStudy (freckles):

\[y=-\sin(x)+Ax+B \\ \text{ we found } A=-2 \\ \text{ and } B=\pi+1 \\ \text{ plug \in and done }\]

OpenStudy (anonymous):

holy crap i found it !!! its D AAAAAHH this is ssoo cool !!!

OpenStudy (freckles):

lol you're funny

OpenStudy (anonymous):

im such a pro !!!!

OpenStudy (anonymous):

okay maybe nmot .... but still !!!

OpenStudy (anonymous):

thanks for the help , i really appreciate it!

OpenStudy (freckles):

np